OpenStudy (anonymous):

Use the quadratic formula and a calculator to solve the following equation: 1.69p2 + 5.98p + 5.29 = 0

OpenStudy (anonymous):

use quadratic equation\[(-b \pm \sqrt{b ^{2}-4ac})\div(2a)\] wher a = 1.63 b=5.98 c=5.29

OpenStudy (anonymous):

The solution will be exact if the decimal numbers are converted to fractions. Solve the following for p:\[\frac{169}{100}p^2 +\frac{598}{100}p + \frac{529}{100} \text{ = } 0\]\[p=-\frac{23}{13}=-1.76923 \]